Determine whether the vectors u and v are parallel, orthogonal, or neither.

u = <6, -2>, v = <8, 24>

help me please
1
Expert's answer
2014-05-15T11:39:10-0400
Vectors are perpendicular if their dot product is zero v*w=06*8-2*24=0They are perpendicular